what volume of 0.676 mol/l sulfuric acid can be neutralized 27 mL of 0.385 moles/liters barium hydroxide
1
Expert's answer
2020-05-12T12:38:06-0400

Ba(OH)2 + SO3 = BaSO4 + H2O.

27 x 0.385 / 1000 = 0,010395.

0,010395/0.676= 0,0153772189.